OpenStudy (anonymous):

HELP PLS solve system of equations 9x+4y = -13 8x+5y=0

OpenStudy (roadjester):

are you familiar with matricies?

OpenStudy (anonymous):

no :(

OpenStudy (roadjester):

What you want to do is get rid of one of the variables

OpenStudy (roadjester):

I'd recommend using the second equation since the constant is 0

OpenStudy (a1234):

9x+4y = -13 8x+5y=0 I'll use elimination: 45x + 20y = -65 32x + 20y = 0 13x = -65 x = -5 9(-5) + 4y = -13 -45 + 4y = -13 4y = 32 y = 8 (-5, 8)

OpenStudy (roadjester):

@a1234 is correct. Just ignore me. =)

OpenStudy (anonymous):

how did you get 45x + 20y = -65 32x + 20y = 0 ????

OpenStudy (jdoe0001):

using elimination as shown by a1234 -> \(\large \begin{array}{llll} 9x+4y=-13&{\color{red}{ \times 5}}\implies & \quad 45x\cancel{+20y}=-65\\ 8x+5y=0&{\color{red}{ \times -4}}\implies &-32x\cancel{-20y}=\quad 0\\ \hline\\ &&13x\qquad \qquad =-65 \end{array}\)